evaluate ∫ 1 / 1 – cos 2x dx

evaluate
∫  1 / 1 – cos 2x dx

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

I = \int \frac{1}{1-cos2x}dx
I = \int \frac{1}{2sin^2x}dx
I = \int \frac{cosec^2x}{2}dx
I = \frac{-cotx}{2} + c